OpenStudy (anonymous):

Find the value of 15 - 8 ÷ 2. 3.5 4 11

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

well this is algebra

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know what PEMDAS is?

OpenStudy (ghostedly):

Answer is 11.

OpenStudy (anonymous):

no i mean i havent sense last year i started working on geometry so im really rusty.

OpenStudy (anonymous):

bro you cant give direct answers

OpenStudy (ghostedly):

I just did.

OpenStudy (ghostedly):

Don't start a problem with me, like that other abdullah guy.

OpenStudy (anonymous):

anyway we would want want to divide first as that comes first in PEMDAS

OpenStudy (anonymous):

so 15-8/2 would be?

OpenStudy (anonymous):

i dont have a calculator but ill try my best in my head...

OpenStudy (anonymous):

no just do 8/2

OpenStudy (anonymous):

11

OpenStudy (anonymous):

8/2 is 4

OpenStudy (ghostedly):

15 - 8/2 = 15/1 - 8/2 = 15/1 + -8/2 = 15/1 +-4/1 = 15 + -4/1 = 11 There my explanation.

OpenStudy (anonymous):

because

OpenStudy (anonymous):

but nice try :D

OpenStudy (anonymous):

so its 4

OpenStudy (anonymous):

8 devide by s is 4

OpenStudy (anonymous):

15-4 is what we have now

OpenStudy (anonymous):

4 -15 =11

OpenStudy (anonymous):

so the answer is 11 as ghostedly said

OpenStudy (anonymous):

feel free to tag me in any other questions you might have

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

no medal ?:(

OpenStudy (ghostedly):

She gave it to me, but I gave you one.

OpenStudy (anonymous):

ty

OpenStudy (ghostedly):

Np.

OpenStudy (ghostedly):

Eh, she took mine away but good job.

OpenStudy (anonymous):

so should i give mr ghostedly his mmetal back

OpenStudy (anonymous):

gave you medal :)

OpenStudy (anonymous):

nvm now we all have one